Suppose that TapDance, Inc.s, capital structure features 70 percent equity, 30 percent debt, and that its before-tax cost of debt is 10 percent, while its cost of equity is 15 percent. Assume the appropriate weighted average tax rate is 34 percent What will be TapDances WACC? (Round your answer to 2 decimal places.)
Answer #1

After tax cost of debt=10(1-tax rate)

=10(1-0.34)=6.6%

WACC=Respective costs*Respective weight

=(0.7*15)+(0.3*6.6)

=12.48%